# Augmenting Multi-Agent Communication with State Delta Trajectory

### v1

@@ Abstract
Multi-agent techniques such as role playing or multi-turn debates have been shown to be effective in improving the performance of large language models (LLMs) in downstream tasks. Despite their differences in workflows, existing LLM-based multi-agent systems mostly use natural language for agent communication. While this is appealing for its simplicity and interpretability, it also introduces inevitable information loss as one model must down sample its continuous state vectors to concrete tokens before transferring them to the other model. Such losses are particularly significant when the information to transfer is not simple facts, but reasoning logics or abstractive thoughts. To tackle this problem, we propose a new communication protocol that transfers both natural language tokens and token-wise state transition trajectory from one agent to another. Particularly, compared to the actual state value, we find that the sequence of state changes in LLMs after generating each token can better reflect the information hidden behind the inference process, so we propose a State Delta Encoding (SDE) method to represent state transition trajectories. The experimental results show that multi-agent systems with SDE achieve SOTA performance compared to other communication protocols, particularly in tasks that involve complex reasoning. This shows the potential of communication augmentation for LLM-based multi-agent systems. [^1] We have open-sourced all the code and data in https://github.com/LittleDinoC/StateDelta/.
@@ 1 Introduction
Multi-agent systems based on Large Language Models (LLMs) have demonstrated remarkable performance in solving complex tasks Guo et al. (2024a); Dong et al. (2024); Du et al. (2024). While it is not surprising that combining outputs from different LLMs could improve the system performance Xu et al. (2023); Chu et al. (2024); Xu et al. (2024), studies have shown that building a multi-agent system with a single base LLM can also boost the LLM’s performance Chan et al. (2024); Hong et al. (2024); Du et al. (2024). These systems construct multiple agents from the same LLM, varying their profiles or access to information, which can be seen as another form of the inference scaling law Chen et al. (2024); Qian et al. (2025). Therefore, how to build effective multi-agent frameworks or workflows to improve LLMs in downstream tasks has been widely studied in recent literature.
@@ 2.2 Multi-agent Communication
Another approach Ramesh and Li (2025) attempts to directly transfer hidden states between agents, but is restricted to a unidirectional transfer, where hidden states from a text-reading agent are transferred to an output-generating agent. It does not support dynamic, interactive exchanges typical in multi-agent systems.
@@ 3 Methodology
To minimize the impact on the model’s generation capabilities, we apply SDE to only a few carefully selected layers. The optimal layers for injection depend on the model’s architecture and scale, but once selected, they work well across various downstream tasks, which indicates that the selection is largely task-agnostic. Layer selection is performed via a simple preliminary experiment and remains fixed for each model throughout all subsequent tasks. Details of our selection process are provided in Section 4.5.
@@ 4.5 Layer Selection
We identify suitable injection layers through a simple preliminary experiment. Specifically, we construct an IA task using the 2WikiMultihopQA Ho et al. (2020) dataset, following the settings described in Section 4.1. For each model, we evaluate SDE’s performance when modifying each layer on the first 300 questions. Considering model scales, we select 1, 2, or 3 layers for 7B, 8B, and 14B models, respectively. These selected layers are then used consistently across all experiments. Notably, 2WikiMultihopQA is used only for this selection procedure and not in any main evaluation. Our main results show that optimal layers depend primarily on the model itself, rather than the downstream task. Further analysis on the impact of different layer selections and layer counts is provided in Section 5.2.
@@ 5.1 Main Experiments
Multi-agent systems perform better than single agents directly answering in most cases. In particular, SDE outperforms existing communication protocols (NL and CIPHER) almost all tasks. These improvements suggest that enriching communication with hidden states can indeed enhance the final collaboration performance of multi-agent systems.
@@
Moreover, we compare our method on StrategyQA using the same model Qwen2.5-7B-Instruct, under two different settings: information asymmetry and agent workflows. Our results show that the agent workflow tasks — which requires more complex reasoning — benefits more significantly from our approach. This also suggests that SDE is particularly effective in tasks that involve more complex reasoning processes.
@@ Appendix A Layer Selection
It is important to note that the 2WikiMultihopQA dataset is used only in this layer selection procedure and is excluded from all evaluations. Our main experimental results suggest that optimal injection layers are primarily determined by model architecture and are relatively robust to specific task settings.
@@ B.1 Information Asymmetry (IA) Tasks
Since each agent has different private information, we implement a single agent answering baseline in which each agent independently performs retrieval-augmented generation based solely on its own private passages. We report the higher of the two agents’ total scores as the baseline performance. Prompt D is used in the single agent baseline.
@@
To ensure reproducibility, we use greedy decoding for Single, Natural Language (NL), and SDE methods, and set the temperature to 0 for CIPHER for fair comparison. Each generation is limited to at most 256 tokens.
@@ B.2.1 Multi-agent Debate
To encourage diverse initial responses under the same first-round prompt, we use randomization during generation. For the Single, Natural Language (NL), and SDE methods, we adopt the model’s default generation settings. For CIPHER, we adjust agent’s temperatures based on the number of agents: in an $n$-agent system, the $i$-th agent use a temperature of $\frac{i}{n}\times$ the model’s default temperature. Specifically, in our 2-agent setting, one agent uses half the default temperature and the other uses the default temperature. The default generation settings for each model are listed below:
@@ Appendix C Different Layer Selections
  • Combined-layer modification (Combine Top-$k$ yields more stable performance across different values of $k$. While in some isolated cases, a single-layer modification may outperform the combined version, the latter demonstrates better robustness and generality across tasks.
@@ Appendix D Prompts
Prompt 7: Prompt example for multi-agent systems used in agent workflow tasks constructed from the HotpotQA dataset and the StrategyQA dataset <user> Solve a question answering task with interleaving Thought, Action, Observation steps. Thought can reason about the current situation, and Action can be three types: (1) Search[entity], which searches the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to search. (2) Lookup[keyword], which returns the next sentence containing keyword in the current passage. (3) Finish[answer], which returns the answer and finishes the task. Here are some examples. Question: What is the elevation range for the area that the eastern sector of the Colorado orogeny extends into? Thought 1: I need to search Colorado orogeny, find the area that the eastern sector of the Colorado orogeny extends into, then find the elevation range of the area. Action 1: Search[Colorado orogeny] Observation 1: The Colorado orogeny was an episode of mountain building (an orogeny) in Colorado and surrounding areas. Thought 2: It does not mention the eastern sector. So I need to look up eastern sector. Thought 3: The eastern sector of Colorado orogeny extends into the High Plains. So I need to search High Plains and find its elevation range. Action 3: Search[High Plains] Observation 3: High Plains refers to one of two distinct land regions: Thought 4: I need to instead search High Plains (United States). Action 4: Search[High Plains (United States)] Observation 4: The High Plains are a subregion of the Great Plains. From east to west, the High Plains rise in elevation from around 1,800 to 7,000 ft (550 to 2,130 m).[3] Thought 5: High Plains rise in elevation from around 1,800 to 7,000 ft, so the answer is 1,800 to 7,000 ft. Action 5: Finish[1,800 to 7,000 ft] ……(five examples omitted)…… Question: Were Scott Derrickson and Ed Wood of the same nationality? Thought 1: I need to determine the nationalities of Scott Derrickson and Ed Wood to see if they are the same. I’ll start by searching for Scott Derrickson. Action 1: Search[Scott Derrickson] Observation 1: Scott Derrickson is an American director, screenwriter and producer. He lives in Los Angeles, California. Derrickson is best known for directing numerous horror films, such as T̈he Exorcism of Emily Rose(̈2005), S̈inister(̈2012), and D̈eliver Us From Evil(̈2014), as well as the Marvel Cinematic Universe superhero film D̈octor Strange(̈2016). He was born on July 16, 1966 in Denver, Colorado, United States. He is married to actress and producer Jennifer Derrickson, with whom he has two children. </user> <assistant> Thought 2:

### v2

@@ Abstract
Multi-agent techniques such as role playing or multi-turn debates have been shown to be effective in improving the performance of large language models (LLMs) in downstream tasks. Despite their differences in workflows, existing multi-agent systems constructed from a single base LLM mostly use natural language for agent communication. While this is appealing for its simplicity and interpretability, it also introduces inevitable information loss as one model must down sample its continuous state vectors to discrete tokens before transferring them to the other model. Such losses are particularly significant when the information to transfer is not simple facts, but reasoning logics or abstractive thoughts. To tackle this problem, we propose a new communication protocol that transfers both natural language tokens and token-wise state transition trajectory from one agent to another. Particularly, compared to the actual state value, we find that the sequence of state changes in LLMs after generating each token can better reflect the information hidden behind the inference process. We propose a State Delta Encoding (SDE) method to represent state transition trajectories. The experimental results show that multi-agent systems with SDE achieve SOTA performance compared to other communication protocols, particularly in tasks that involve complex reasoning. [^1] We have open-sourced all the code and data in https://github.com/LittleDinoC/StateDelta/.
@@ 1 Introduction
Multi-agent systems based on Large Language Models (LLMs) have demonstrated remarkable performance in solving complex tasks Guo et al. (2024a); Dong et al. (2024); Du et al. (2024). While it is not surprising that combining outputs from different LLMs could improve system performance Xu et al. (2023); Chu et al. (2024); Xu et al. (2024), studies have shown that building a multi-agent system with a single base LLM can also boost the LLM’s performance Chan et al. (2024); Hong et al. (2024); Du et al. (2024). These systems construct multiple agents from the same LLM, varying their profiles or access to information, which can be seen as another form of the inference scaling law Chen et al. (2024); Qian et al. (2025). Therefore, methods for building effective multi-agent frameworks or workflows to improve LLMs in downstream tasks have been widely studied in recent literature.
@@ 2.2 Multi-agent Communication
Another approach Ramesh and Li (2025) attempts to directly transfer hidden states between agents. In their setup, communication is restricted to a unidirectional transfer, where hidden states from a text-reading agent are passed to an output-generating agent. Moreover, in the multi-agent debate setting, their method requires heterogeneous models to provide distinct activations, whereas ours operates with multiple instances of the same model.
@@ 3 Methodology
To minimize the impact on the model’s generation capabilities, we apply SDE to only a few carefully selected layers. The layers are chosen based on a simple preliminary experiment and remain fixed for each model throughout all subsequent tasks. Our results show that the layers selected in this way consistently yield good results on the downstream tasks we evaluate, without requiring task-specific layer selection. Details of our selection process are provided in Section 4.5.
@@ 4.1 Information Asymmetry (IA) Tasks
Although our passage split cannot strictly guarantee that the agents’ private information is completely disjoint, we design a single-agent baseline to demonstrate that the knowledge available to any single agent alone is insufficient to answer the questions. In this baseline, each agent independently performs retrieval-augmented generation using only its private passages, and we report the better of the two agents’ scores. The performance gap between this single-agent baseline and the multi-agent systems highlights the necessity of information exchange: only by transmitting complementary private knowledge can the agents successfully complete the task.
@@ 4.5 Layer Selection
We identify suitable injection layers through a simple preliminary experiment. Specifically, we construct an IA task using the 2WikiMultihopQA Ho et al. (2020) dataset, following the settings described in Section 4.1. For each model, we evaluate SDE’s performance when modifying each layer on the first 300 questions. Considering model scales, we select 1, 2, or 3 layers for 7B, 8B, and 14B models, respectively. These selected layers are then used consistently across all experiments. Notably, 2WikiMultihopQA is used only for this selection procedure and not in any main evaluation. While the selected layers are not guaranteed to be optimal, our experiments show that they still generalize effectively to other downstream tasks we used without additional tuning. Further analysis on the impact of different layer selections and layer counts is provided in Section 5.2.
@@ 5.1 Main Experiments
Multi-agent systems perform better than single agents directly answering in most cases. In particular, SDE outperforms existing communication protocols (NL and CIPHER) on almost all tasks. These improvements suggest that enriching communication with hidden states can indeed enhance the final collaboration performance of multi-agent systems.
@@
Moreover, we compare our method on StrategyQA using the same model Qwen2.5-7B-Instruct, under two different settings: information asymmetry and agent workflows. Our results show that the agent workflow tasks — which require more complex reasoning — benefit more significantly from our approach. This also suggests that SDE is particularly effective in tasks that involve more complex reasoning processes.
@@ 7 Limitation
Acknowledgments
This research was supported by Tsinghua University Initiative Scientific Research Program.
@@ Appendix A Layer Selection
It is important to note that the 2WikiMultihopQA dataset is used only in this layer selection procedure and is excluded from all evaluations.
@@ B.1 Information Asymmetry (IA) Tasks
We adopt two agents mainly to balance computational efficiency and model performance, since involving more agents substantially increases context length and may harm generation quality. Similarly, we choose 6 passages because in comparable RAG setups a single model typically accesses three documents; in our two-agent setting, this corresponds to three private documents per agent, totaling six.
@@
Since each agent has different private information, we implement a single-agent answering baseline in which each agent independently performs retrieval-augmented generation based solely on its own private passages. We report the higher of the two agents’ total scores as the baseline performance. Prompt D is used in the single-agent baseline.
We intentionally did not evaluate a single agent with access to the union of all retrieved passages from both agents. Our focus is on assessing whether the proposed multi-agent communication protocol can effectively transmit agent-specific private information to improve performance. This single-agent design aligns with that goal: the observed performance gap, where independent agents with only private information underperform compared to communicating agents, demonstrates that the agents have complementary knowledge and that successful information transmission is crucial for task completion.
@@
Each generation is limited to at most 256 tokens. To ensure reproducibility, we use greedy decoding for Single, Natural Language (NL), and SDE methods, and set the temperature to 0 for CIPHER for fair comparison. We note that, for CIPHER, setting temperature to 0 does not make the output equivalent to a one-hot selection. Following the official implementation [^2] https://github.com/chaudatascience/cipher_multiagent_debate/blob/main/models/agent.py#L768C1-L772C1, when temperature is 0, the softmax is computed directly on the raw logits, which is effectively the same as setting the temperature to 1. Thus, the generated CIPHER embedding still captures the full probability distribution over the vocabulary.
@@ B.2.1 Multi-agent Debate
To encourage diverse initial responses under the same first-round prompt, we use randomization during generation. For the Single, Natural Language (NL), and SDE methods, we adopt the model’s default generation settings. In CIPHER, each agent encodes a temperature-scaled version of the model’s logits into its embedding. Agents using the same temperature would produce identical embeddings, leading to no differences in debates. To address this, CIPHER assigns different temperatures to each agent to encourage varied embeddings. In our experiments, we follow the original implementation and set the $i$-th agent’s temperature to $\frac{i}{n}\times T_{\text{default}}$, where $n$ is the total number of agents and $T_{\text{default}}$ is the model’s default temperature. This simple scheme preserves the original CIPHER design—having one agent with a higher temperature and one with a lower temperature—without the need for dataset-specific tuning. The default generation settings for each model are listed below:
@@ Appendix C Different Layer Selections
  • Combined-layer modification (Combine Top-$k$) yields more stable performance across different values of $k$. While in some isolated cases, a single-layer modification may outperform the combined version, the latter demonstrates better robustness and generality across tasks.
@@ Appendix D Prompts
Prompt 7: Prompt example for multi-agent systems used in agent workflow tasks constructed from the HotpotQA dataset and the StrategyQA dataset <user> Solve a question answering task with interleaving Thought, Action, Observation steps. Thought can reason about the current situation, and Action can be three types: (1) Search[entity], which searches the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to search. (2) Lookup[keyword], which returns the next sentence containing keyword in the current passage. (3) Finish[answer], which returns the answer and finishes the task. Here are some examples. Question: What is the elevation range for the area that the eastern sector of the Colorado orogeny extends into? Thought 1: I need to search Colorado orogeny, find the area that the eastern sector of the Colorado orogeny extends into, then find the elevation range of the area. Action 1: Search[Colorado orogeny] Observation 1: The Colorado orogeny was an episode of mountain building (an orogeny) in Colorado and surrounding areas. Thought 2: It does not mention the eastern sector. So I need to look up eastern sector. Thought 3: The eastern sector of Colorado orogeny extends into the High Plains. So I need to search High Plains and find its elevation range. Action 3: Search[High Plains] Observation 3: High Plains refers to one of two distinct land regions: Thought 4: I need to instead search High Plains (United States). Action 4: Search[High Plains (United States)] Observation 4: The High Plains are a subregion of the Great Plains. From east to west, the High Plains rise in elevation from around 1,800 to 7,000 ft (550 to 2,130 m).[3] Thought 5: High Plains rise in elevation from around 1,800 to 7,000 ft, so the answer is 1,800 to 7,000 ft. Action 5: Finish[1,800 to 7,000 ft] ……(five examples omitted)…… Question: Were Scott Derrickson and Ed Wood of the same nationality? Thought 1: I need to determine the nationalities of Scott Derrickson and Ed Wood to see if they are the same. I’$l$ start by searching for Scott Derrickson. Action 1: Search[Scott Derrickson] Observation 1: Scott Derrickson is an American director, screenwriter and producer. He lives in Los Angeles, California. Derrickson is best known for directing numerous horror films, such as T̈he Exorcism of Emily Rose(̈2005), S̈inister(̈2012), and D̈eliver Us From Evil(̈2014), as well as the Marvel Cinematic Universe superhero film D̈octor Strange(̈2016). He was born on July 16, 1966 in Denver, Colorado, United States. He is married to actress and producer Jennifer Derrickson, with whom he has two children. </user> <assistant> Thought 2: